Harry spends $62 on four CDs two DVDs. the next day he spends $57 on two CDs and 3 DVDs. how much does one CD and one DVD cost?
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Harry spends $62 on four CDs two DVDs. the next day he spends $57 on two CDs and 3 DVDs. how much does one CD and one DVD cost?
--------
4C + 2D = 62
2C + 3D = 57
--------------------
Modify for elimination:
4C + 2D = 62
4C + 6D = 114
-----------------------
Subtract and solve for "D"::
4D = 52
D = 12 (Cost of a DVD)
---
Solve for "C"::
2C + 3D = 57
24 + 3D = 57
3D = 33
D = 11 (Cost of a CD)
----------
Cheers,
Stan H.
-----------

Answer by MathTherapy(10810)


Harry spends $62 on four CDs two DVDs. the next day he spends $57 on two CDs and 3 DVDs. how much does one CD and one DVD cost?